Skip to Search

 
#

Support Services Division Meeting
October 29, 2007

Library West Admin Conference Room, 557
Present: Bill Covey, Kelley Cunningham, Erich Kesse, Barbara Oliver, Rob Roberts, Will Chaney
Absent: Debra Fetzer, Brian Keith
Recorded by: Amanda Humphries
Business Services:
  1. Payroll closes a day early this week due to the Homecoming Holiday. Please tell your staff to have their hours in by 9am Wednesday, October 31.
  2. Reference has requested a second Suggestions Box for Library West 3rd Floor. Bill and Barbara will discuss this at another time.
  3. Business Services is prepared to initiate the second half of TEAMS Reclassification but is awaiting additional information.
  4. The Pre-Bid Security System meeting scheduled for Tuesday, October 30 will be post-poned. Apparently bids have not yet gone out.
Facilities:
  1. The Smathers Library Room 450 asbestos tile will cost $1900 to remove. Rob will work with Barbara to determine how to appropriate funds for this expense.
  2. The terrazzo guy will be able to give an estimate for the refinishing of East after he finishes business with the Reitz Union. Bill suggested gathering estimates for both terrazzo and marble.
  3. Pam Cenzer met with the decorator for the MSL repairs. They have prepared a package to present to Dean Russell.
  4. Facilities is receiving requests from Departments for additional safety measures (i.e. safety vaults, automatic door lock over-ride software, etc.) These requests will be denied.
  5. The Security Slide-Card conversions will be a complete conversion for all branches.
  6. Bill will meet with Purchasing regarding how to handle the Lutron software copy issue.
  7. The supply inventory in Facilities is not dwindling quickly enough. Barbara will send out a reminder to the Supplies List-Serv reminding contacts to check with Peter before ordering. Additionally, all supplies taken from the Facilities inventory will no longer count against Departmental Supplies Budgets.
Systems:
  1. The Greenstone Product machines are up and running. The issue turned out to be an outside multiple access problem.
  2. The SQL Server problem appears to be a full log. Logan is looking into it.
  3. It was decided that Aleph print demons will not be replaced, but out-dated multi-function machines should be.
  4. Will will start the Mp3 move from the server tomorrow.
  5. Systems will give extra Cisco switches to CNS and retain approximately six.
  6. The Acquisitions request for a 30-inch flat panel was denied.
Digital Library Center:
  1. The DLC Storage drives have arrived and will be installed in time for the DLC to back up portable storage drives in anticipation of meeting the Florida Newspaper vendor.
  2. The three department meeting (DLC, Preservation & Special Collections) will be revived via a Wiki. The DLC is working on setting it up. They will employ word-mapping technology to display overlapping projects.
Bill Covey reports:
  1. The Directors will meet today to discuss the OPS Budget, after which OPS allocations will be finalized.
  2. The $500,000 extra is to be used for enhancement, most likely in materials.
  3. The DSR funds have come back in one lump sum. The Directors are working on sorting out which funds will go where.
  4. This Friday is the Homecoming Holiday.
  5. The one-time $1,000 bonuses will be reflected in the November 9th paycheck.
Meeting adjourned.

 

University of Florida Home Page